Circular No. 8215

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2003 T1 (TRITTON) P. Holvorcem, Campinas, Brazil, has reported that the co-addition of

three 45-s unfiltered CCD images of a fast-moving object found by C. Juels, Fountain Hills, AZ, with a 0.12-m f/5 refractor and a 0.5-m f/4.8 reflector (discovery observation below) show a coma of diameter 2′ and a hint of a 1′.5 tail at p.a. ∼ 257^{o}. Following posting on the NEO Confirmation Page, additional CCD observers noted the object’s cometary appearance, including R. Trentman (Louisburg, KS, 0.75-m reflector; mag 13.1 and very faint evidence of a tail ≈ 10′′ long in p.a. ≈ 280^{o} on Oct. 7.4 UT), D. T. Durig (Sewanee, TN, 0.30-m f/5.86 reflector; teardrop-shaped coma of mag 10.1 with a tail at least 2′–2′.5 long in p.a. ∼ 285^{o} on Oct. 7.4), and J. Young (Table Mountain, CA, 0.6-m reflector; 36′′ coma elongated to 48′′, with a 3′ tail in p.a. 289^{o} with a very straight and extremely thin jet of length ∼ 1′.5 in its center on Oct. 7.5).

2003 UT  α_{2000}  δ_{2000}  Mag.
Oct. 6.44284  9^{h}08^{m}52.^{s}72  +18^{o}58′15′′.9  12.4

Following a suggestion by S. Hönig (Dossenheim, Germany) from orbital computations by M. Meyer (Kelkheim, Germany), B. G. Marsden (Smithsonian Astrophysical Observatory) has shown that this comet is identical to the lost comet 1978d = 1977 XIII = D/1978 C2 (Tritton), which was observed for only a month (cf. IAUC 3175, 3186, 3194, 3198). The available astrometry, including Sept. 22 prediscovery observations, and the following orbital elements by Marsden appear on MPEC 2003-T37.

Epoch = 1977 Oct. 24.0 TT

T = 1977 Oct. 28.4469 TT  ω = 147.^{o}2879
e = 0.587351  Ω = 300.9159 }2000.0
q = 1.435688 AU  i = 7.0812

a = 3.479198 AU n^{o} = 0.1518747 P = 6.490 years

Epoch = 2003 Oct. 8.0 TT

T = 2003 Sept.24.3077 TT  ω = 147.^{o}4233
e = 0.589828  Ω = 300.7297 }2000.0
q = 1.421098 AU  i = 7.1157

a = 3.464642 AU n^{o} = 0.1528328 P = 6.449 years

2003 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green